# Graph Positional Autoencoders as Self-supervised Learners

### v1

@@ 4. Proposed Framework: GraphPAE
In this section, we introduce GraphPAE, a novel GAE designed to reconstruct both the node features and positions. There are three key components in GraphPAE: corruption, encoder, and decoder, as illustrated in Figure 2(a).
@@ 4.1. Data Corruption
The masking-predicting paradigm has become a basic component of autoencoders. A well-designed masking strategy can prevent information leakage and enhance model efficiency (He et al., 2022). Existing GAEs commonly apply the random masking strategy to the discrete nodes and edges. However, node positions, such as eigenvectors, are continuous by nature and are not suitable for the random masking approach. Therefore, we adopt different corruption strategies for these two modalities.
Mask Feature. We follow GraphMAE (Hou et al., 2022) and replace the masked node features with a learnable vector. Specifically, we randomly sample a subset of nodes $\widetilde{\mathcal{V}}\subset\mathcal{V}$ and reset their features. The corrupted feature matrix $\widetilde{\mathbf{X}}$ is defined as:
@@
Offset Position. Inspired by the recent advances in molecular representation learning (Zhou et al., 2023), we propose to add random offsets to node positions. Specifically, we use the top-$k$ eigenvectors of graph Laplacian $\textbf{U}\in\mathbb{R}^{N\times K}$ as the initialization of node positions. Similar to the feature corruption, we define the corruption of node positions as:
@@ 4.2. GraphPAE Encoder
where $l\in\{0,1,2,...,L\}$ indicates the layer of encoder, and $\mathcal{N}_{i}$ is the neighbors of node $v_{i}$. The first layer of the encoder, e.g., $f_{\text{enc}}^{(0)}$, is designed to align the dimensions of node and position representations. Specifically, it first lifts the scalar relative distance $\mathbf{P}_{i,j}\in\mathbb{R}$ into a vector representation $\mathbf{P}_{i,j}^{(0)}\in\mathbb{R}^{d_{h}}$ through a series of Gaussian RBF kernels
@@
Without loss of generality, we omit edge features in the formulation. Regarding normalization function in Figure 2(b), GAT uses Softmax function, whereas GatedGCN applies degree normalization.
@@ 4.3. GraphPAE Decoder
where $\alpha$ is the hyperparameter. The pseudocode of GDEM is presented in Algorithm 1.
@@ 5. Experiments
In this section, we conduct extensive experiments, including unsupervised node classification, unsupervised graph prediction, and transfer learning on large-scale molecule graphs to verify the effectiveness of GraphPAE. Moreover, we perform ablation studies on position reconstruction and dual-path design. Finally, we analyze the influence of the position reconstruction loss.
@@ 5.1. Node Classification
Dataset. We evaluate the performance of GraphPAE on 6 representative heterophilic graphs: BlogCatalog (Meng et al., 2019), Chameleon, Squirrel, Actor (Pei et al., 2020), arXiv-year (Hu et al., 2020b), and Penn94 (Traud et al., 2012). Specifically, arXiv-year and Penn94 are large-scale graphs (¿ 40,000) to evaluate the scalability of the methods. For all the datasets, we use the public splits.
@@
5.2. Graph Prediction
Results. Table 3 summarizes the results on graph-level tasks. GraphPAE consistently achieves superior performance across both regression and classification benchmarks, highlighting its ability to learn high-quality graph representations. Notably, GraphPAE shows competitive performance against contrastive learning methods and even exhibits notable improvements compared against the spectral-based method $\text{Sp}^{2}\text{GCL}$ on molfreesolv, molbace, and molclintox. In addition, GraphPAE also outperforms other GAEs across most datasets. We attribute these gains to the proposed position corruption-reconstruction strategy, which improves the encoder’s ability to identify crucial substructures for downstream tasks.
@@ 5.3. Transfer Learning
Settings. We conduct transfer learning experiments on molecule property prediction to evaluate the generalization ability of GraphPAE. Specifically, we follow the setting of (Liu et al., 2024a), which first pre-trains the encoder on 2 million molecules sampled from ZINC15 (Sterling and Irwin, 2015), and then fine-tunes on QM9 (Wu et al., 2018) to predict the quantum chemistry properties. We compare GraphPAE with the state-of-the-art molecule graph pre-training model, i.e., GraphCL (You et al., 2020), GraphMAE (Hou et al., 2022), Mole-BERT (Xia et al., 2023), and SimSGT (Liu et al., 2024a). We use a five-layer GatedGCN with 300 hidden units. Upon finishing pre-training, a two-layer MLP is attached after the graph representations for property prediction. During the fine-tuning protocol, the encoder and attached MLP are trained together on labeled data of downstream tasks. Consistent with (Liu et al., 2024a), we divided QM9 into train/validation/test sets with 80%/10%/10% by scaffold splits. We use Adam optimizer and run 5 times to report average MAE and standard deviation. The results are reported in Table 4.
Results. We observe that GraphPAE demonstrates robust generalization across all prediction targets compared to state-of-the-art models. Among the baselines, Mole-BERT and SimSGT are specially designed for molecular graph pre-training, incorporating customized reconstruction objectives tailored to the unique characteristics of molecular graphs. Specifically, Mole-BERT pretrains a discrete codebook within the subgraph of nodes, inherently capturing both node features and local structural patterns. During reconstruction, the model replaces the raw features with codebook entries as predictive targets. SimSGT employs a tokenizer to encode feature and local structure information, using the tokenized outputs as reconstruction targets. Despite these tailored strategies, GraphPAE consistently outperforms these models. The superiority is largely attributed to the effective positional encodings obtained by position corruption-reconstruction, enabling the graph representations to capture crucial substructures for prediction targets such as cycles.
@@ 5.4. Ablation Studies
Effectiveness of Position Reconstruction. To achieve position corruption-reconstruction, GraphPAE inevitably introduces positional encodings into the encoder. Firstly, we verify that the superior performance is attributed not only to the introduction of positional encoding but also to the position reconstruction. We conducted two sets of comparative experiments in Table 5: one comparing Exp a with Exp b, and another comparing Exp c with Exp d. We have the following summaries: (1) In Exp a, both features and positions are corrupted, but only the features are reconstructed. In Exp b, both features and positions are reconstructed. We observe that Exp b performs better than Exp a, demonstrating the effectiveness of position reconstruction. (2) In Exp c, we remove both position corruption and reconstruction from GraphPAE. Comparing Exp d against Exp c, we find that the encoder incorporating position reconstruction consistently improves performance over that solely with feature reconstruction.
@@ 5.5. Parameters Analysis
Influence of the loss weight $\alpha$ in GraphPAE. Figure 3(a) and 3(b) presents the hyperparameter analysis of $\alpha$ to further examine the influence of the positional reconstruction loss $\mathcal{L}_{pos}$. We summarize the key observations as follows: (1) As $\alpha$ increases within a certain threshold, the performance improves progressively, indicating that position reconstruction contributes positively to the quality of learned representations. (2) However, excessively large values of $\alpha$ lead to performance degeneration. This suggests assigning a high weight to position reconstruction may cause the encoder to overemphasize positional information at the expense of overall representation learning. (3) While the optimal value of $\alpha$ varies slightly across datasets, we observe that the best-performing range typically lies within [1e-3, 1e-2, 1e-1], making it easy to search in practice. Additional hyperparameter details are provided in Appendix A.4.
Influence of the number of eigenvectors $K$ in GraphPAE. We further analyze the effect of the number of eigenvectors $K$ on performance using the Squirrel and Chameleon datasets, as shown in Figure 3(c) and 3(d). As $K$ increases, performance generally improves due to the model’s ability to capture a wider range of frequency information. However, once $K$ surpasses a certain threshold, the performance gains become marginal. Therefore, to balance effectiveness and computational efficiency, we adopt a moderate value of $K$, typically in the range of [50-100], which provides a good trade-off in practice.
@@ Acknowledgements.
Appendix A Experimental Details
A.1. Spectral Analysis of Masking Strategies
Masking Strategies for Node Features and Edges. We conduct the experiments in the Squirrel datasets with 20% masking ratio. For node feature masking, given the node feature matrix $\mathbf{X}$, we randomly mask 20% of the node in the dataset and set the feature vector to zero. The remaining 80% of nodes retain their original feature. The corrupted feature matrix is defined as $\widetilde{\mathbf{X}}$. For edge masking, given the original graph structure $\mathbf{A}$, we randomly remove 20% of the edges from the graph while preserving the node features. This process results in a structurally corrupted graph with a new adjacency matrix $\widetilde{\mathbf{A}}$.
@@ A.1. Spectral Analysis of Masking Strategies
For the feature masking case, we calculate $\mathbf{L}$ by substituting $\mathbf{A}$ into Equation 18. Then we decompose $\mathbf{L}$ and get $\mathbf{U}$. By substituting $\mathbf{U}$, $\mathbf{X}$ and $\mathbf{U}$, $\widetilde{\mathbf{X}}$ in Equation 19-22, we can get the frequency magnitude of original and corrupted graphs. The comparison of differences is shown in Figure 1(a).
For edge masking case, we calculate $\mathbf{L}$ and $\widetilde{\mathbf{L}}$ by substituting $\mathbf{A}$ and $\widetilde{\mathbf{A}}$ into Equation 18. Then we decompose $\mathbf{L}$ and $\widetilde{\mathbf{L}}$, and get $\mathbf{U}$ and $\widetilde{\mathbf{U}}$. By substituting $\mathbf{U}$, $\widetilde{\mathbf{U}}$ and $\mathbf{X}$ in Equation (19) - (22), we can also get the frequency magnitude of original and corrupted graphs as shown in Figure 1(b).
A.2. Additional Experiment Results.
We further evaluate GraphPAE on homophilic graphs Facebook (Rozemberczki et al., 2021) and Wiki (Mernyei and Cangea, 2020), as well as a large-scale graph ogbn-Products (Hu et al., 2020b). Notably, ogbn-Products contains 2,449,029 nodes and 61,859,140 edges, making it a suitable benchmark to demonstrate the scalability of GraphPAE. The experimental results are summarized in Table 7, and we make the following observations: (1) GraphPAE consistently achieves competitive performance across these different benchmarks. (2) The performance gains on homophilic graphs are less pronounced compared to those on heterophilic graphs. We conjecture this is because, in homophilic graphs, a small portion of very low-frequency information is sufficient.
Methods Facebook Wiki ogbg-Products
BGRL 89.71±0.35 79.02±0.13 78.59±0.02
CCA-SSG 89.45±0.60 78.85±0.32 75.27±0.05
GraphMAE 89.54±0.36 78.94±0.48 78.89±0.01
GraphMAE2 88.49±0.43 78.84±0.44 81.59±0.02
GraphPAE 91.46±0.23 79.32±0.29 79.10±0.02
A.3. Statistics of Datasets
We benchmark GraphPAE across various tasks, including node classification, graph prediction, and transfer learning. Specifically, we conduct experiments on six node classification datasets and seven graph-level prediction datasets and use ZINC-2M and QM9 for transfer learning. For datasets in node classification and graph prediction, we adopt public data splits. In transfer learning, we follow (Liu et al., 2024a) and divide QM9 into train/valid/test sets with 80%/10%/10% by scaffold split. Detailed statistics of the node- and graph-level datasets are presented in Table 8 and Table 9, respectively.
A.4. Hyperparamters
Hyperparameter details of node classification and graph prediction are reported in Table 10 and Table 11, respectively. $r$ denotes the mask ratio and $\alpha$ is the loss weight of $\mathcal{L}_{pos}$. dp and $\text{dp}_{edge}$ represent the dropout of node features and edge features, respectively. For transfer learning, we pre-train GraphPAE on 2 million molecules sampled from ZINC15 with 100 epochs. The mask ratio is set to 0.35, and the loss weight for $\mathcal{L}_{pos}$ is 0.01. Both node and edge dropout are set to 0.0, and the number of eigenvectors for positional encoding is 6. During fine-tuning, the encoder and the attached MLP are jointly trained on the QM9 dataset using an initial learning rate of 0.001 and no weight decay. We apply a dropout rate of 0.1 to both node features and edge features within GNN layers.
A.5. Complexity Analysis
During training, GraphPAE introduces an additional $\mathcal{O\left(\mathcal{E}\right)}$ complexity per layer for PE computation and an extra cost for relative node distance reconstruction. Since we only reconstruct corrupted node pairs, the complexity remains much lower than $\mathcal{O}\left(\mathcal{E}\right)$. During inference, the extra overhead also comes from PE computation.
Overall, GraphPAE introduces an additional $\mathcal{O}(\mathcal{E})$ complexity in training. However, this slight overhead enables the encoder to capture more diverse frequency information.

### v2

@@ 4. Proposed Framework: GraphPAE
In this section, we introduce GraphPAE, a novel GAE designed to reconstruct both node features and positions. There are three key components in GraphPAE: corruption, encoder, and decoder, as illustrated in Figure 2(a).
@@ 4.1. Data Corruption
The corruption-reconstruction paradigm has become a basic component of autoencoders. A well-designed masking strategy can prevent information leakage and enhance model efficiency (He et al., 2022). Existing GAEs commonly apply the random masking strategy to the discrete nodes and edges. However, node positions, such as eigenvectors, are continuous by nature and are not suitable for the random masking approach. Therefore, we adopt different corruption strategies for these two modalities.
Feature Masking. We follow GraphMAE (Hou et al., 2022) and replace the masked node features with a learnable vector. Specifically, we randomly sample a subset of nodes $\widetilde{\mathcal{V}}\subset\mathcal{V}$ and reset their features. The corrupted feature matrix $\widetilde{\mathbf{X}}$ is defined as:
@@
Position Offsetting. Inspired by the recent advances in molecular representation learning (Zhou et al., 2023), we propose to add random offsets to node positions. Specifically, we use the top-$k$ eigenvectors of graph Laplacian $\textbf{U}\in\mathbb{R}^{N\times K}$ as the initialization of node positions. Similar to the feature corruption, we define the corruption of node positions as:
@@ 4.2. GraphPAE Encoder
where $l\in\{0,1,2,...,L\}$ indicates the layer of encoder, and $\mathcal{N}_{i}$ is the neighbors of node $v_{i}$. The first layer of the encoder, e.g., $f_{\text{enc}}^{(0)}$, is designed to align the dimensions of node features and position representations. Specifically, it first lifts the scalar relative distance $\mathbf{P}_{i,j}\in\mathbb{R}$ into a vector representation $\mathbf{P}_{i,j}^{(0)}\in\mathbb{R}^{d_{h}}$ through a series of Gaussian RBF kernels
@@
Without loss of generality, we omit edge features in the formulation. Regarding the normalization function in Figure 2(b), GAT uses the Softmax function, and GatedGCN applies degree normalization.
@@ 4.3. GraphPAE Decoder
where $\alpha$ is the hyperparameter. The pseudocode of GraphPAE is presented in Algorithm 1.
@@ 5. Experiments
In this section, we conduct extensive experiments, including node classification, graph prediction, and transfer learning on large-scale molecule graphs to verify the effectiveness of GraphPAE. Moreover, we perform ablation studies on position reconstruction and dual-path design. Finally, we analyze the influence of the loss of weight and the number of eigenvectors.
@@ 5.1. Node Classification
Dataset. We evaluate the performance of GraphPAE on 6 representative heterophilic graphs: BlogCatalog (Meng et al., 2019), Chameleon, Squirrel, Actor (Pei et al., 2020), arXiv-year (Hu et al., 2020b), and Penn94 (Traud et al., 2012). Specifically, arXiv-year and Penn94 are large-scale graphs (¿ 40,000) to evaluate the scalability of the methods. As these datasets place greater emphasis on high-frequency information, they are well suited to evaluate GraphPAE’s ability to capture diverse frequency components.
@@
5.2. Graph Prediction
Results. Table 3 summarizes the results of graph-level tasks. GraphPAE consistently achieves superior performance across both regression and classification benchmarks, highlighting its ability to learn high-quality graph representations. Notably, GraphPAE shows competitive performance against contrastive learning methods and even exhibits notable improvements compared against the spectral-based method $\text{Sp}^{2}\text{GCL}$ on molfreesolv, molbace, and molclintox. In addition, GraphPAE also outperforms other GAEs across most datasets. We attribute these gains to the proposed position corruption-reconstruction strategy, which improves the encoder’s ability to identify crucial substructures for downstream tasks.
@@ 5.3. Transfer Learning
Settings. We conduct transfer learning experiments on molecule property prediction to evaluate the generalization ability of GraphPAE. Specifically, we follow the setting of (Liu et al., 2024a), which first pre-trains the encoder on 2 million molecules sampled from ZINC15 (Sterling and Irwin, 2015), and then fine-tunes on QM9 (Wu et al., 2018) to predict the quantum chemistry properties. We compare GraphPAE with the state-of-the-art molecule graph pre-training model, i.e., GraphCL (You et al., 2020), GraphMAE (Hou et al., 2022), Mole-BERT (Xia et al., 2023), and SimSGT (Liu et al., 2024a). We use a five-layer GatedGCN with 300 hidden units. Upon finishing pre-training, a two-layer MLP is attached after the graph representations for property prediction. During the fine-tuning protocol, the encoder and attached MLP are trained together on labeled data of downstream tasks. Consistent with (Liu et al., 2024a), we divided QM9 into train/validation/test sets with 80%/10%/10% by scaffold splits. We use Adam optimizer and run it 5 times to report average MAE and standard deviation.
Results. From Table 4, we observe that GraphPAE demonstrates robust generalization across all prediction targets compared to state-of-the-art models. Among the baselines, Mole-BERT and SimSGT are specially designed for molecular graph pretraining, incorporating customized reconstruction objectives tailored to the characteristics of molecular graphs. Specifically, Mole-BERT pretrains a discrete codebook within the subgraph of nodes, inherently capturing both features and local structural patterns. During reconstruction, the model replaces the raw features with codebook entries as predictive targets. SimSGT employs a tokenizer to encode feature and local structure information, using tokenized output as reconstruction targets. Despite these tailored strategies, GraphPAE consistently outperforms these models. The superiority is largely attributed to the effective positional encodings obtained by position corruption-reconstruction, enabling the graph representations to capture crucial substructures for prediction targets such as cycles.
@@ 5.4. Ablation Studies
Effectiveness of Position Reconstruction. To achieve position corruption-reconstruction, GraphPAE inevitably introduces positional encodings into the encoder. Firstly, we verify that the superior performance is attributed not only to the introduction of positional encodings but also to the position reconstruction. We conducted two sets of comparative experiments in Table 5: one comparing Exp a with Exp b, and another comparing Exp c with Exp d. We have the following summaries: (1) In Exp a, both features and positions are corrupted, but only the features are reconstructed. In Exp b, both features and positions are reconstructed. We observe that Exp b performs better than Exp a, demonstrating the effectiveness of position reconstruction. (2) In Exp c, we remove both position corruption and reconstruction from GraphPAE. Comparing Exp d against Exp c, we find that the encoder incorporating position reconstruction consistently improves performance compared to feature reconstruction alone.
@@ 5.5. Parameters Analysis
Influence of the loss weight $\alpha$ in GraphPAE. Figure 3(a) and 3(b) present the hyperparameter analysis of $\alpha$ to further examine the influence of the positional reconstruction loss $\mathcal{L}_{pos}$. We summarize the key observations as follows: (1) As $\alpha$ increases within a certain threshold, the performance improves progressively, indicating that position reconstruction contributes positively to the quality of learned representations. (2) However, excessively large values of $\alpha$ lead to performance degeneration. This suggests assigning a high weight to position reconstruction may cause the encoder to overemphasize positional information at the expense of overall representation learning. (3) While the optimal value of $\alpha$ varies slightly across datasets, we observe that the best-performing range typically lies within {1e-3, 1e-2, 1e-1}, making it easy to search in practice. Additional hyperparameter details are provided in Appendix B.2.
Influence of the number of eigenvectors $K$ in GraphPAE. We further analyze the effect of the number of eigenvectors $K$ on performance using the Squirrel and Chameleon datasets, as shown in Figure 3(c) and 3(d). As $K$ increases, performance generally improves due to the model’s ability to capture a wider range of frequency information. However, once $K$ surpasses a certain threshold, the performance gains become marginal. Therefore, to balance effectiveness and computational efficiency, we adopt a moderate value of $K$, typically in the range of [50, 100], which provides a good trade-off in practice.
@@ Acknowledgements.
Appendices
Appendix A Spectral Analysis of Masking Strategies
Masking Strategies for Node Features and Edges. We conduct the experiments in the Squirrel datasets with a 20% masking ratio. For node feature masking, given the node feature matrix $\mathbf{X}$, we randomly mask 20% of the node in the dataset and set the feature vector to zero. The remaining 80% of nodes retain their original feature. The corrupted feature matrix is defined as $\widetilde{\mathbf{X}}$. For edge masking, given the original graph structure $\mathbf{A}$, we randomly remove 20% of the edges from the graph while preserving the node features. This process results in a structurally corrupted graph with a new adjacency matrix $\widetilde{\mathbf{A}}$.
@@ Appendix A Spectral Analysis of Masking Strategies
For feature masking case, we calculate $\mathbf{L}$ by substituting $\mathbf{A}$ into Equation 18. Then we decompose $\mathbf{L}$ and get $\mathbf{U}$. By substituting $\mathbf{U}$, $\mathbf{X}$, and $\widetilde{\mathbf{X}}$ in Equation (19)-(22), we can get the frequency magnitude of original and corrupted graphs. The comparison of differences is shown in Figure 1(a). For edge masking case, we calculate $\mathbf{L}$ and $\widetilde{\mathbf{L}}$ by substituting $\mathbf{A}$ and $\widetilde{\mathbf{A}}$ into Equation 18. Then we decompose $\mathbf{L}$ and $\widetilde{\mathbf{L}}$, and get $\mathbf{U}$ and $\widetilde{\mathbf{U}}$. By substituting $\mathbf{U}$, $\widetilde{\mathbf{U}}$, and $\mathbf{X}$ in Equation (19) - (22), we can also get the frequency magnitude of original and corrupted graphs as shown in Figure 1(b).
@@
Appendix B Experimental Details
B.1. Statistics of Datasets
We benchmark GraphPAE across various tasks, including node classification, graph prediction, and transfer learning. Specifically, we conduct experiments on six node classification datasets and seven graph-level prediction datasets and select ZINC15 and QM9 for transfer learning. For datasets in node classification and graph prediction, we utilize public data splits. In transfer learning, we follow (Liu et al., 2024a) and divide QM9 into train/valid/test sets with 80%/10%/10% by scaffold split. Detailed statistics of the node- and graph-level datasets are presented in Table 7 and Table 8, respectively.
B.2. Hyperparameters
Hyperparameter details of node classification and graph prediction are reported in Table 9 and Table 10. $r$ denotes the mask ratio and $\alpha$ is the loss weight of $\mathcal{L}_{pos}$. dp and $\text{dp}_{edge}$ represent the dropout of node features and edge features, respectively. For transfer learning, we pre-train GraphPAE on 2 million molecules sampled from ZINC15 with 100 epochs. The mask ratio is set to 0.35, and the loss weight for $\mathcal{L}_{pos}$ is 0.01. Both node and edge dropout are set to 0.0, and the number of eigenvectors for positional encoding is 6. During fine-tuning, the encoder and the attached MLP are jointly trained on the QM9 dataset using an initial learning rate of 0.001 and no weight decay. We apply a dropout rate of 0.1 to both node features and edge features within GNN layers.
B.3. Additional Experiment Results.
We further evaluate GraphPAE on homophilic graphs, including Facebook (Rozemberczki et al., 2021) and Wiki (Mernyei and Cangea, 2020), as well as a large-scale graph ogbn-products (Hu et al., 2020b). Notably, ogbn-products contains 2,449,029 nodes and 61,859,140 edges, making it a suitable benchmark to demonstrate the scalability of GraphPAE. The experimental results are summarized in Table 11, and we make the following observations: (1) GraphPAE consistently achieves competitive performance across these different benchmarks. (2) The performance gains on homophilic graphs are less pronounced compared to those on heterophilic graphs. We conjecture this is because, in homophilic graphs, a small portion of very low-frequency information is sufficient.
Appendix C Complexity Analysis
During training, GraphPAE introduces an additional $\mathcal{O\left(\mathcal{E}\right)}$ complexity per layer for PE computation and an extra cost for relative node distance reconstruction. Since we only reconstruct corrupted node pairs, the complexity remains much lower than $\mathcal{O}\left(\mathcal{E}\right)$. However, this slight overhead enables the encoder to capture more diverse frequency information.
@@ Appendix C Complexity Analysis
Methods Facebook Wiki ogbg-products
BGRL 89.71±0.35 79.02±0.13 78.59±0.02
CCA-SSG 89.45±0.60 78.85±0.32 75.27±0.05
GraphMAE 89.54±0.36 78.94±0.48 78.89±0.01
GraphMAE2 88.49±0.43 78.84±0.44 81.59±0.02
GraphPAE 91.46±0.23 79.32±0.29 79.10±0.02